OpenStudy (anonymous):

Find the sum of the infinite geometric series: 1.6 + 1.2 + 0.9 + ... Can someone help me solve this?

OpenStudy (anonymous):

sum = a (1-r) here r = 1.2 / 1.6 = 0.75 = 1.6 /( 1 - 0.75) = 6.4

OpenStudy (anonymous):

a/(1-r) = sum to infinity where |r| < 1

OpenStudy (anonymous):

oooh, thanks a lot jimmyrep! ^_^ I get it now

OpenStudy (anonymous):

good